graph slope intercept method 10x -2y =4
y=mx + b
-2y = -10x + 4
-2y/-2= -10x/-2 + 4/-2
y= 5x +-2
am I on the right track
: graph slope intercept method 10x -2y =4
y=mx + b
-2y = -10x + 4
-2y/-2= -10x/-2 + 4/-2
y= 5x +-2
am I on the right track

Answer by checkley77(3848)

10x -2y =4
y=mx + b
-2y = -10x + 4
-2y/-2= -10x/-2 + 4/-2
y= 5x +-2 [drop the + before the -2]
am I on the right track?
YES.
 graph( 300, 200, -6, 5, -10, 10, 5x -2) (graph 300x200 pixels, x from -6 to 5, y from -10 to 10, 5x -2).
SLOPE (m)=5
Y INTERCEPT (b)=-2